OpenStudy (anonymous):

determine the exact value of the trig ratios: cos(13pi/4) cot(11pi/2) sec(5pi/3)

OpenStudy (jamesj):

Remember that ALL trig functions have periodicity 2pi e.g., sin(x) = sin(x + 2pi), cos(x) = cos(x + 2pi), ...

OpenStudy (jamesj):

So for example, cot(11pi/2) = cot(6pi - pi/2) = cot(-pi/2)

OpenStudy (anonymous):

thank you!! i forgot about that!

OpenStudy (jamesj):

This isn't the only trick you'll need, but it will help you get these expressions in terms you'll be able to handle more easily. For example, for the cot example above: cot(-pi/2) = cos(-pi/2)/sin(-pi/2) = 0/-1 = 0

OpenStudy (jamesj):

sure thing